OpenStudy (solomonzelman):

I am not very good at this stuff, so trying to grasp it with more practice....

OpenStudy (solomonzelman):

\({\huge{\bbox[5pt, lightyellow ,border:2px solid black ]{ \color{lightyellow}{\frac{\frac{\color{black}{\rm Hey~~guys,~~\color{blue}{c}\color{green}{o}\color{blue}{n}\color{green}{g}\color{blue}{r}\color{green}{a}\color{blue}{t}\color{green}{u}\color{blue}{l}\color{green}{a}\color{blue}{t}\color{green}{i}\color{blue}{o}\color{green}{n}!~~~~~~~~~~~~~~~~~~~~~~~} }{ \color{black}{\rm I~~feel~~this~~is~~a~~very~~big~~achievement!~~ } } }{ \frac{\color{black}{\rm I~~am~~really~~happy~~for~~you~~guys,~~you~~} }{ \color{black}{\rm are~~just~~awesome!~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ ~~} } }} }}}\)

OpenStudy (solomonzelman):

oh, it actually comes out handy!!

OpenStudy (solomonzelman):

\({\huge{\triangle \triangle\triangle\triangle\triangle\triangle\triangle\triangle\triangle\triangle\triangle\triangle\triangle \\} {\bbox[5pt, lightyellow ,border:2px solid black ]{ \color{lightyellow}{\frac{\frac{\color{black}{\rm Hey~~guys,~~\color{blue}{c}\color{green}{o}\color{blue}{n}\color{green}{g}\color{blue}{r}\color{green}{a}\color{blue}{t}\color{green}{u}\color{blue}{l}\color{green}{a}\color{blue}{t}\color{green}{i}\color{blue}{o}\color{green}{n}!~~~~~~~~~~~~~~~~~~~~~~~} }{ \color{black}{\rm I~~feel~~this~~is~~a~~very~~big~~achievement!~~ } } }{ \frac{\color{black}{\rm I~~am~~really~~happy~~for~~you~~guys,~~you~~} }{ \color{black}{\rm are~~just~~awesome!~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ ~~} } }} }}}\)

OpenStudy (solomonzelman):

I am not done, I will practice more. MODERATORS, please don't close it

Nnesha (nnesha):

i'm not goood at all :

OpenStudy (confluxepic):

Maybe add some more color.

OpenStudy (solomonzelman):

well, just inserting more colors is not really a technique enhancing...

OpenStudy (solomonzelman):

No, I want to border the sides of this box with squares too, but it seems as thoguh as that is not a possibility

OpenStudy (confluxepic):

@Nnesha did it.

OpenStudy (solomonzelman):

bordered the box from sides as well as top and bottom the box with squares?

OpenStudy (confluxepic):

Yes.

OpenStudy (confluxepic):

Kind of.

OpenStudy (solomonzelman):

I want:like that